# Homological stratification and descent

### 04 Mar 2026 v2

@@ 4. Homological stratification via cosupport
Consider the following statements for a rigidly-compactly generated tt-category $\mathscr{T}$:
@@
A rigidly-compactly generated tt-category $\mathscr{T}$ is h-stratified if and only if the h-LGP holds for $\mathscr{T}$ and
@@
For a rigidly-compactly generated tt-category $\mathscr{T}$, the following are equivalent:
@@ 7. Weakly descendable functors and descent
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a family of geometric functors between rigidly-compactly generated tt-categories, indexed on a (not necessarily finite) set $I$. We refer to $(f_{i}^{*})_{i\in I}$ as a geometric family.
@@
in $\mathscr{T}$. The situation studied in loc. cit. is the special case when $I$ is a singleton and $f^{*}$ is given by base-change along a commutative algebra.
@@
$(a)\Rightarrow(c)$: If $t_{1}\in\operatorname{Locid}\langle t_{2}\rangle$ then $f_{i}^{*}(t_{1})\in\operatorname{Locid}\langle f_{i}^{*}(t_{2})\rangle$ for all $i\in I$, since $f_{i}^{*}$ are tt-functors, so it remains to prove the reverse implication in $(c)$. Tensoring (7.1) with $t_{1}$ and using the projection formula gives
@@
In light of the above corollary and its analogue [BCH+23b, Corollary 14.24], we regard weak descendability as the appropriate condition for studying descent properties in rigidly-compactly generated tt-geometry.
@@
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category. By definition, the family of localizations $\{f_{\mathscr{P}}^{*}\colon\mathscr{T}\to\mathscr{T}_{\mathscr{P}}\}_{\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})}$ is weakly descendable if and only if
@@
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category. If the homological local-to-global principle holds then the family of localizations
@@
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category whose spectrum is weakly noetherian. If the local-to-global principle holds then the family of localizations
@@
Suppose $\mathscr{T}$ admits enough tt-fields, say $\{f_{\mathscr{B}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{B}}\}_{\mathscr{B}}$. The following are equivalent:
@@
If $\mathscr{T}$ has the h-detection property then $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle(f_{\mathscr{B}})_{*}\mathbb{1}\rangle$ for every $\mathscr{B}$ by 3.5, which implies the desired equivalence. Certainly h-LGP implies h-detection. To finish the proof, we claim that $(a)$ also implies h-detection. Indeed, since tt-fields are h-stratified (3.4), we have $(\varphi^{h}_{\mathscr{B}})^{-1}(\operatorname{Supp}^{h}(t))=\operatorname{Supp}^{h}(f_{\mathscr{B}}^{*}(t))$ for every $\mathscr{B}$ and every object $t\in\mathscr{T}$ by 6.8. Therefore, the h-detection property is equivalent to the family $\{f_{\mathscr{B}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{B}}\}_{\mathscr{B}}$ being jointly conservative. The claim thus follows from 7.9. ∎
@@
8. Comparison between h-stratification and tt-stratification
A theory of stratification for tt-categories is developed in [BHS23d] under the assumption that the Balmer spectrum is weakly noetherian. Our next goal is to relate this notion of stratification (based on the Balmer spectrum and the Balmer–Favi support) with the homological stratification developed in this paper (based on the homological spectrum and the homological support).
@@ 8. Comparison between h-stratification and tt-stratification
We will say that a rigidly-compactly generated tt-category $\mathscr{T}$ satisfies the Nerves of Steel conjecture ([BAL20b, Remark 5.15]) if the surjective comparison map
@@
The question of whether every rigidly-compactly generated tt-category satisfies the Nerves of Steel conjecture is an open problem, but it is known to be true in many examples of interest; see [BAL20b, Section 5], for example. For the state of the art, see [HYS24].
@@
We will say that a rigidly-compactly generated tt-category whose spectrum $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian is tt-stratified if it is stratified in the sense of [BHS23d].
@@
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category with $\operatorname{Spc}(\mathscr{T}^{c})$ weakly noetherian. The following are equivalent:
@@
$(a)\Rightarrow(b)$: By [BHS23c, Theorem 4.7] if $\mathscr{T}$ is tt-stratified, then the Nerves of Steel conjecture holds for $\mathscr{T}$, so it remains to show that $\mathscr{T}$ is h-stratified. It was proved in [BHS23c, Lemma 3.7] that $\operatorname{Supp}(E_{\mathscr{B}})=\{\pi(\mathscr{B})\}$ for any $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$. Hence $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle g_{\pi(\mathscr{B})}\rangle$ when $\mathscr{T}$ is tt-stratified. Therefore,
@@
for any $t\in\mathscr{T}$. This coincides with $\operatorname{Locid}\langle E_{\mathscr{B}}\mid\mathscr{B}\in\operatorname{Supp}^{h}(t)\rangle$ since [BHS23c, Theorem 4.7] establishes that $\pi^{-1}(\operatorname{Supp}(t))=\operatorname{Supp}^{h}(t)$ when $\mathscr{T}$ is tt-stratified. This establishes that $\mathscr{T}$ is h-stratified by 3.1.
@@
commutes. By hypothesis, the horizontal and vertical maps are injective (and thus bijective), hence so is the diagonal map; that is, $\mathscr{T}$ is tt-stratified. ∎
@@
If the Nerves of Steel conjecture holds for $\mathscr{T}$ and $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian then $\mathscr{T}$ is tt-stratified if and only if it is h-stratified.
@@
Corollary 8.8.
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a weakly descendable geometric family. Assume that $\mathscr{S}_{i}$ is tt-stratified for each $i\in I$ and that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian. If the Nerves of Steel conjecture holds for $\mathscr{T}$ then $\mathscr{T}$ is tt-stratified.
@@
For each $i\in I$, the $t$-category $\mathscr{S}_{i}$ is tt-stratified by assumption, hence h-stratified by 8.6. Descent for h-stratification as established in 7.22 then implies that $\mathscr{T}$ is h-stratified. Therefore, 8.6 implies that $\mathscr{T}$ is tt-stratified, due to our assumption that $\mathscr{T}$ satisfies Nerves of Steel. ∎
@@ 9. Descent for the Nerves of Steel conjecture
In order to apply 8.8 to descend tt-stratification, we desire methods for establishing the Nerves of Steel conjecture. This is the topic of the present section.
@@
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category. Suppose there exists a tt-field $f_{\mathscr{P}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{P}}$ detecting $\mathscr{P}$ for each prime $\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})$, i.e., $\mathscr{P}$ is in the image of $\operatorname{Spc}(f_{\mathscr{P}}^{*})$. The following are equivalent:
@@
Recall from [BC21, Lemma 2.2] that every tt-field induces a homological residue field. More precisely, for each $\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})$, there exists a homological prime $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$ lying over $\mathscr{P}$ and a commutative diagram
@@
where the functor $\mkern 1.5mu\overline{\mkern-1.5muF\mkern-1.5mu}\mkern 1.5mu$ is faithful; see [BC21, (2.3)]. It follows from this diagram and the fact that the tt-field $\mathscr{F}_{\mathscr{P}}$ is phantomless that a morphism in $\mathscr{T}$ is killed by $f_{\mathscr{P}}^{*}$ if and only if it is killed by the homological residue field $\overline{\mathbb{h}}_{\mathscr{B}}$. Thus, the family $(f_{\mathscr{P}}^{*})_{\mathscr{P}\in\mathscr{\operatorname{Spc}}(\mathscr{T}^{c})}$ jointly detects $\otimes$-nilpotence if and only if the corresponding family of homological residue fields (also indexed on $\operatorname{Spc}(\mathscr{T}^{c})$) jointly detects $\otimes$-nilpotence. The latter is equivalent to the Nerves of Steel conjecture by [BAL20b, Corollary 4.7 and Theorem 5.4]. This establishes the equivalence of $(a)$ and $(b)$. On the other hand, contemplating the commutative diagram
@@ 10. Applications to tt-stratification
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a weakly descendable geometric family with each $\mathscr{S}_{i}$ tt-stratified. Suppose that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian. Assume additionally one of the following:
@@
Then $\mathscr{T}$ is tt-stratified.
@@
In the situation of 10.1(c)(c), if the indexing set $I$ is finite then $\operatorname{Spc}(\mathscr{T}^{c})$ is automatically noetherian and hence the hypothesis that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian is superfluous.
@@
Let $\mathscr{C}$ be a rigidly-compactly generated symmetric monoidal stable $\infty$-category and let $A\in\operatorname{CAlg}(\mathscr{C}^{c})$ be a descendable dualizable commutative algebra in $\mathscr{C}$. If $\operatorname{Mod}_{\mathscr{C}}(A)$ is tt-stratified and has a noetherian spectrum, then $\mathscr{C}$ is also tt-stratified and also has a noetherian spectrum.
@@
Since $A$ is descendable, the base-change functor $A\otimes-\colon\mathscr{C}\to\operatorname{Mod}_{\mathscr{C}}(A)$ is conservative and hence induces a surjection $\operatorname{Spc}(\operatorname{Mod}_{\mathscr{C}}(A)^{c})\to\operatorname{Spc}(\mathscr{C}^{c})$; see [BCH+24b, Theorem 1.4] or [BCH+24a, Proposition 7.8]. Hence $\operatorname{Spc}(\mathscr{C}^{c})$ is also noetherian. We may then invoke 10.1(c)(c). ∎
@@
The previous corollary strengthens the results of [BCH+24a, Section 12] by removing the hypothesis that the descendable algebra $A$ be separable of finite tt-degree.
@@
A rigidly-compactly generated tt-category with weakly noetherian spectrum is said to be cohomologically stratified if it is tt-stratified and the comparison map ([BAL10])
@@
in which the three indicated vertical maps are surjections. The top and middle horizontal maps are also surjections by 7.9. The bottom map is a bijection by assumption, hence $\rho_{\mathscr{T}}$ is surjective as well. Since the tt-categories $\mathscr{S}_{i}$ are cohomologically stratified, the left vertical composite is a bijection, which then implies that all maps in this diagram are bijections. The result thus follows from 8.8. ∎

### 18 Mar 2026 v3

@@ 4. Homological stratification via cosupport
Consider the following statements for a rigidly-compactly generated $t$-category $\mathscr{T}$:
@@
A rigidly-compactly generated $t$-category $\mathscr{T}$ is h-stratified if and only if the h-LGP holds for $\mathscr{T}$ and
@@
For a rigidly-compactly generated $t$-category $\mathscr{T}$, the following are equivalent:
@@ 7. Weakly descendable functors and descent
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a family of geometric functors between rigidly-compactly generated $t$-categories, indexed on a (not necessarily finite) set II. We refer to $(f_{i}^{*})_{i\in I}$ as a geometric family.
@@
in $\mathscr{T}$. The situation studied in loc. cit. is the special case when II is a singleton and $f^{*}$ is given by base-change along a commutative algebra.
@@
$(a)\Rightarrow(c)$: If $t_{1}\in\operatorname{Locid}\langle t_{2}\rangle$ then $f_{i}^{*}(t_{1})\in\operatorname{Locid}\langle f_{i}^{*}(t_{2})\rangle$ for all $i\in I$, since $f_{i}^{*}$ are $t$-functors, so it remains to prove the reverse implication in $(c)$. Tensoring (7.1) with $t_{1}$ and using the projection formula gives
@@
In light of the above corollary and its analogue [BCH+23b, Corollary 14.24], we regard weak descendability as the appropriate condition for studying descent properties in rigidly-compactly generated $t$-geometry.
@@
Let $\mathscr{T}$ be a rigidly-compactly generated $t$-category. By definition, the family of localizations $\{f_{\mathscr{P}}^{*}\colon\mathscr{T}\to\mathscr{T}_{\mathscr{P}}\}_{\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})}$ is weakly descendable if and only if
@@
Let $\mathscr{T}$ be a rigidly-compactly generated $t$-category. If the homological local-to-global principle holds then the family of localizations
@@
Let $\mathscr{T}$ be a rigidly-compactly generated $t$-category whose spectrum is weakly noetherian. If the local-to-global principle holds then the family of localizations
@@
Suppose $\mathscr{T}$ admits enough $t$-fields, say $\{f_{\mathscr{B}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{B}}\}_{\mathscr{B}}$. The following are equivalent:
@@
If $\mathscr{T}$ has the h-detection property then $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle(f_{\mathscr{B}})_{*}\mathbb{1}\rangle$ for every $\mathscr{B}$ by 3.5, which implies the desired equivalence. Certainly h-LGP implies h-detection. To finish the proof, we claim that $(a)$ also implies h-detection. Indeed, since $t$-fields are h-stratified (3.4), we have $(\varphi^{h}_{\mathscr{B}})^{-1}(\operatorname{Supp}^{h}(t))=\operatorname{Supp}^{h}(f_{\mathscr{B}}^{*}(t))$ for every $\mathscr{B}$ and every object $t\in\mathscr{T}$ by 6.8. Therefore, the h-detection property is equivalent to the family $\{f_{\mathscr{B}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{B}}\}_{\mathscr{B}}$ being jointly conservative. The claim thus follows from 7.9. ∎
@@
8. Comparison between h-stratification and $t$-stratification
A theory of stratification for $t$-categories is developed in [BHS23d] under the assumption that the Balmer spectrum is weakly noetherian. Our next goal is to relate this notion of stratification (based on the Balmer spectrum and the Balmer–Favi support) with the homological stratification developed in this paper (based on the homological spectrum and the homological support).
@@ 8. Comparison between h-stratification and $t$-stratification
We will say that a rigidly-compactly generated $t$-category $\mathscr{T}$ satisfies the Nerves of Steel conjecture ([BAL20b, Remark 5.15]) if the surjective comparison map
@@
The question of whether every rigidly-compactly generated $t$-category satisfies the Nerves of Steel conjecture is an open problem, but it is known to be true in many examples of interest; see [BAL20b, Section 5], for example. For the state of the art, see [HYS24].
@@
We will say that a rigidly-compactly generated $t$-category whose spectrum $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian is $t$-stratified if it is stratified in the sense of [BHS23d].
@@
Let $\mathscr{T}$ be a rigidly-compactly generated $t$-category with $\operatorname{Spc}(\mathscr{T}^{c})$ weakly noetherian. The following are equivalent:
@@
$(a)\Rightarrow(b)$: By [BHS23c, Theorem 4.7] if $\mathscr{T}$ is $t$-stratified, then the Nerves of Steel conjecture holds for $\mathscr{T}$, so it remains to show that $\mathscr{T}$ is h-stratified. It was proved in [BHS23c, Lemma 3.7] that $\operatorname{Supp}(E_{\mathscr{B}})=\{\pi(\mathscr{B})\}$ for any $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$. Hence $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle g_{\pi(\mathscr{B})}\rangle$ when $\mathscr{T}$ is $t$-stratified. Therefore,
@@
for any $t\in\mathscr{T}$. This coincides with $\operatorname{Locid}\langle E_{\mathscr{B}}\mid\mathscr{B}\in\operatorname{Supp}^{h}(t)\rangle$ since [BHS23c, Theorem 4.7] establishes that $\pi^{-1}(\operatorname{Supp}(t))=\operatorname{Supp}^{h}(t)$ when $\mathscr{T}$ is $t$-stratified. This establishes that $\mathscr{T}$ is h-stratified by 3.1.
@@
commutes. By hypothesis, the horizontal and vertical maps are injective (and thus bijective), hence so is the diagonal map; that is, $\mathscr{T}$ is $t$-stratified. ∎
@@
If the Nerves of Steel conjecture holds for $\mathscr{T}$ and $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian then $\mathscr{T}$ is $t$-stratified if and only if it is h-stratified.
@@
Corollary $8.$
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a weakly descendable geometric family. Assume that $\mathscr{S}_{i}$ is $t$-stratified for each $i\in I$ and that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian. If the Nerves of Steel conjecture holds for $\mathscr{T}$ then $\mathscr{T}$ is $t$-stratified.
@@
For each $i\in I$, the $t$-category $\mathscr{S}_{i}$ is $t$-stratified by assumption, hence h-stratified by 8.6. Descent for h-stratification as established in 7.22 then implies that $\mathscr{T}$ is h-stratified. Therefore, 8.6 implies that $\mathscr{T}$ is $t$-stratified, due to our assumption that $\mathscr{T}$ satisfies Nerves of Steel. ∎
@@ 9. Descent for the Nerves of Steel conjecture
In order to apply 8.8 to descend $t$-stratification, we desire methods for establishing the Nerves of Steel conjecture. This is the topic of the present section.
@@
Let $\mathscr{T}$ be a rigidly-compactly generated $t$-category. Suppose there exists a $t$-field $f_{\mathscr{P}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{P}}$ detecting $\mathscr{P}$ for each prime $\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})$, i.e., $\mathscr{P}$ is in the image of $\operatorname{Spc}(f_{\mathscr{P}}^{*})$. The following are equivalent:
@@
Recall from [BC21, Lemma 2.2] that every $t$-field induces a homological residue field. More precisely, for each $\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})$, there exists a homological prime $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$ lying over $\mathscr{P}$ and a commutative diagram
@@
where the functor $\mkern 1.5mu\overline{\mkern-1.5muF\mkern-1.5mu}\mkern 1.5mu$ is faithful; see [BC21, (2.3)]. It follows from this diagram and the fact that the $t$-field $\mathscr{F}_{\mathscr{P}}$ is phantomless that a morphism in $\mathscr{T}$ is killed by $f_{\mathscr{P}}^{*}$ if and only if it is killed by the homological residue field $\overline{\mathbb{h}}_{\mathscr{B}}$. Thus, the family $(f_{\mathscr{P}}^{*})_{\mathscr{P}\in\mathscr{\operatorname{Spc}}(\mathscr{T}^{c})}$ jointly detects $\otimes$-nilpotence if and only if the corresponding family of homological residue fields (also indexed on $\operatorname{Spc}(\mathscr{T}^{c})$) jointly detects $\otimes$-nilpotence. The latter is equivalent to the Nerves of Steel conjecture by [BAL20b, Corollary 4.7 and Theorem 5.4]. This establishes the equivalence of $(a)$ and $(b)$. On the other hand, contemplating the commutative diagram
@@ 10. Applications to $t$-stratification
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a weakly descendable geometric family with each $\mathscr{S}_{i}$ $t$-stratified. Suppose that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian. Assume additionally one of the following:
@@
Then $\mathscr{T}$ is $t$-stratified.
@@
In the situation of 10.1(c)(c), if the indexing set II is finite then $\operatorname{Spc}(\mathscr{T}^{c})$ is automatically noetherian and hence the hypothesis that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian is superfluous.
@@
Let $\mathscr{C}$ be a rigidly-compactly generated symmetric monoidal stable $\infty$-category and let $A\in\operatorname{CAlg}(\mathscr{C}^{c})$ be a descendable dualizable commutative algebra in $\mathscr{C}$. If $\operatorname{Mod}_{\mathscr{C}}(A)$ is $t$-stratified and has a noetherian spectrum, then $\mathscr{C}$ is also $t$-stratified and also has a noetherian spectrum.
@@
Since AA is descendable, the base-change functor $A\otimes-\colon\mathscr{C}\to\operatorname{Mod}_{\mathscr{C}}(A)$ is conservative and hence induces a surjection $\operatorname{Spc}(\operatorname{Mod}_{\mathscr{C}}(A)^{c})\to\operatorname{Spc}(\mathscr{C}^{c})$; see [BCH+24b, Theorem 1.4] or [BCH+24a, Proposition 7.8]. Hence $\operatorname{Spc}(\mathscr{C}^{c})$ is also noetherian. We may then invoke 10.1(c)(c). ∎
@@
The previous corollary strengthens the results of [BCH+24a, Section 12] by removing the hypothesis that the descendable algebra AA be separable of finite $t$-degree.
@@
A rigidly-compactly generated $t$-category with weakly noetherian spectrum is said to be cohomologically stratified if it is $t$-stratified and the comparison map ([BAL10])
@@
in which the three indicated vertical maps are surjections. The top and middle horizontal maps are also surjections by 7.9. The bottom map is a bijection by assumption, hence $\rho_{\mathscr{T}}$ is surjective as well. Since the $t$-categories $\mathscr{S}_{i}$ are cohomologically stratified, the left vertical composite is a bijection, which then implies that all maps in this diagram are bijections. The result thus follows from $8.$ ∎